# AN ACCELERATED PRIMAL DUAL ALGORITHM WITH BACKTRACKING FOR DECENTRALIZED CONSTRAINED OPTIMIZATION

### v1

@@ 6. Convergence guarantees for D-APDB and the Proof of Main Results
where we used $\eta^{k}(\alpha_{i}^{k}+\beta_{i}^{k})=(c_{\alpha}+c_{\beta})/\tau_{i}^{k}$ for the first inequality, and $\delta\leq 1-(c_{\alpha}+c_{\beta}+c_{\varsigma})$ in the second inequality. Therefore, for all $i\in\mathcal{N}$, since $\tau_{i}^{k}\leq\bar{\tau}_{i}$, using $\|x_{i}^{k+1}-x_{i}^{k}\|^{2}\leq 4D_{i}^{2}$ and $\|\theta_{i}^{k+1}-\theta_{i}^{k}\|^{2}\leq 4B_{i}^{2}$ together with $\delta\in(0,1)$, we get
$\bar{P}_{i}^{k}\leq-\frac{\delta}{2\tau_{i}^{k}}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}-\frac{\delta}{2\sigma_{i}^{k}}\|\theta_{i}^{k+1}-\theta_{i}^{k}\|^{2}+\Xi_{i},\quad\forall~k\in\mathcal{I},$ (6.4)
Thus, the desired results immediately follows from the bound on the cardinality of $\mathcal{I}$ given in Lemma 6.2 and the fact that $t_{k}\leq t_{0}=1$ for all $k\geq 0$. ∎
@@
Under the premise of Theorem 6.3. Let $\{x_{i}^{k},\theta_{i}^{k}\}_{k\geq 0}$ be the D-APDB iterate sequence for $i\in\mathcal{N}$, and $\{s_{i}^{k}\}_{k\geq 0}$ denote the D-APDB auxiliary sequence with $s_{i}^{0}=0$ for $i\in\mathcal{N}$, and define $\{\lambda^{k}\}_{k\geq 0}$ such that $\lambda^{k}=A\mathbf{s}^{k}$ with $\mathbf{s}^{k}=[s_{i}^{k}]_{i\in\mathcal{N}}$ for all $k\geq 0$. For any $\mathbf{x}\in\mathop{\bf dom}\phi$, $\theta\in\mathcal{K}^{*}\cap\mathcal{B}$ and $\lambda\in\mathbb{R}^{n|\mathcal{E}|}$, let $\mathbf{z}\triangleq(\mathbf{x},\mathbf{y})$ with $\mathbf{y}\triangleq(\theta,\lambda)$. Then, for any $K\geq 1$, it holds that
@@
holds for any $K\geq 1$, where $T_{K}=\sum_{k=0}^{K-1}t_{k}$ and $(\bar{\mathbf{x}}^{K},\bar{\mathbf{y}}^{K})\triangleq\frac{1}{T_{K}}\sum_{k=0}^{K-1}t_{k}(\mathbf{x}^{k},\mathbf{y}^{k})$. Therefore, Lemma 4.3 and Theorem 6.3 together imply that
@@
Using the same arguments we employed while deriving eq. 4.4 and eq. 4.13, we can further lower bound the left hand side of eq. 6.7 as below:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma^{K}(\mathbf{z})}$ $\geq\mathcal{L}(\bar{\mathbf{x}}^{K},\mathbf{y})-\mathcal{L}(\mathbf{x},\bar{\mathbf{y}}^{K})+\frac{1}{2\gamma^{K-1}}\frac{t_{K-1}}{T_{K}}\|\lambda-\lambda^{K}\|^{2}+\frac{t_{K-1}}{2T_{K}}\sum_{i\in\mathcal{N}}\varsigma_{i}^{K}\|x_{i}^{K}-x_{i}^{K-1}\|^{2}$ (6.8)
@@
Let $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathcal{K}^{*}$ denote an arbitrary primal-dual optimal pair for (1.1) as defined in Assumption 4 such that $\|\theta_{i}^{*}\|\leq B_{i}/2$ for $i\in\mathcal{N}$, and $\lambda^{*}\in\mathbb{R}^{n|\mathcal{E}|}$ be an optimal dual variable corresponding to the consensus constraint $A\mathbf{x}=0$ in (2.1). Under the premise of Theorem 6.3, initialized from arbitrary $x_{i}^{0}\in\mathop{\bf dom}\phi_{i}$ and $\theta_{i}^{0}=0$ for all $i\in\mathcal{N}$, the ergodic iterate sequences of D-APDB, i.e., $\{(\bar{x}_{i}^{k},\bar{\theta}_{i}^{k})\}_{k\geq 0}$ for all $i\in\mathcal{N}$, satisfy
@@
$\mathcal{L}(\bar{\mathbf{x}}^{K},\theta^{*},\lambda^{*})-\mathcal{L}(\mathbf{x}^{*},\bar{\theta}^{K},\bar{\lambda}^{K})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Gamma}(\mathbf{x}^{*},\mathbf{y}^{*})}/T_{K}.$ (6.11)
@@
$\tilde{w}_{i}\triangleq g_{i}(\bar{x}_{i}^{K})\in\mathbb{R}^{m_{i}},\quad\forall~i\in\mathcal{N}.$ (6.12)
Since $\mathcal{K}_{i}$ is a closed convex cone, it induces a decomposition on $\mathbb{R}^{m_{i}}$:
@@
which also satisfy $\left\langle\tilde{w}_{i}^{(1)},~\tilde{w}_{i}^{(2)}\right\rangle=0$ for $i\in\mathcal{N}$. Thus, for all $i\in\mathcal{N}$, we have
$\|\tilde{w}_{i}^{(2)}\|=\|\mathcal{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\mathcal{K}_{i}}}(\tilde{w}_{i})-\tilde{w}_{i}\|\triangleq d_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\mathcal{K}_{i}}}(\tilde{w}_{i}).$ (6.14)
Next, define $\tilde{\theta}=[\tilde{\theta}_{i}]_{i\in\mathcal{N}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathcal{K}^{*}\cap\mathcal{B}}$ such that
$\tilde{\theta}_{i}\triangleq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\|\theta_{i}^{*}\|}\cdot\frac{\tilde{w}_{i}^{(2)}}{\|\tilde{w}_{i}^{(2)}\|}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{K}_{i}^{*}\cap\mathcal{B}_{i}},\quad\forall~i\in\mathcal{N}.$ (6.15)
@@
Similarly, define $\tilde{\lambda}\triangleq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\|\lambda^{*}\|}\cdot\frac{A\bar{\mathbf{x}}^{K}}{\|A\bar{\mathbf{x}}^{K}\|}$; hence, from the definition we get
@@
Note that $\mathcal{L}(\mathbf{x}^{*},\bar{\theta}^{K},\bar{\lambda}^{K})\leq\mathcal{L}(\mathbf{x}^{*},\theta^{*},\lambda^{*})=\varphi^{*}$ since $(\mathbf{x}^{*},\theta^{*},\lambda^{*})$ is a saddle point of $\mathcal{L}$, and that $\tilde{\theta}\in\mathcal{K}^{*}\cap\mathcal{B}$; therefore, together with (6.16) and (6.17), we get
@@
Moreover, since $\mathbf{x}^{*}\in\mathop{\bf dom}\phi$ and $\tilde{\theta}\in\mathcal{K}^{*}\cap\mathcal{B}$, eq. 6.5 also implies that
@@
Moreover, using the saddle point property of $(\mathbf{x}^{*},\theta^{*},\lambda^{*})$ one more time, we also have $\mathcal{L}(\bar{\mathbf{x}}^{K},{\theta}^{*},{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},\theta^{*},\lambda^{*})\geq 0$; therefore, we get
@@
Recall that $g_{i}(\bar{x}_{i}^{K})=\tilde{w}_{i}$ such that (6.13) holds for $i\in\mathcal{N}$. Hence, we immediately get
$\langle\theta_{i}^{*},g_{i}(\bar{x}_{i}^{K})\rangle=\langle\theta_{i}^{*},~\tilde{w}_{i}-\tilde{w}_{i}^{(1)}+\tilde{w}_{i}^{(1)}\rangle\leq\langle\theta_{i}^{*},~\tilde{w}_{i}-\tilde{w}_{i}^{(1)}\rangle\leq\|\theta_{i}^{*}\|d_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\mathcal{K}_{i}}}(g_{i}(\bar{x}_{i}^{K}));$ (6.23)
therefore, together with eq. 6.22, we get
@@
Combining eq. 6.21 and eq. 6.24 immediately yields the desired result. ∎
@@
Under the premise of corollary 6.5, there exists $(x^{*},\theta^{*})$ a primal-dual optimal solution to (1.1) such that the actual primal-dual iterate sequence $\{(x_{i}^{k},\theta_{i}^{k})\}_{k\geq 0}$ converges to $(x^{*},\theta_{i}^{*})$ for all $i\in\mathcal{N}$, i.e., $\lim_{k\to\infty}x_{i}^{k}=x^{*}$ and $\lim_{k\to\infty}\theta_{i}^{k}=\theta_{i}^{*}$ for $i\in\mathcal{N}$.
@@
[…]
$\geq\frac{\hat{t}}{2}\sum_{i\in\mathcal{N}}\Big(\frac{\delta}{\bar{\tau}_{i}}\|x_{i}^{k}-\hat{x}_{i}\|^{2}+\frac{1}{\zeta_{i}\bar{\tau}_{i}}\|\theta_{i}^{k}-\hat{\theta}_{i}\|^{2}\Big)+\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\hat{t}\rho}{c_{\alpha}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|\mathcal{E}|}~\min_{i\in\mathcal{N}}\{\hat{\tau}_{i}\}~\|\lambda^{k}-\hat{\lambda}\|^{2}>0,$
$\frac{1}{\gamma^{k}}-\sum_{i\in\mathcal{N}}d_{i}\frac{\eta^{k}}{\varsigma_{i}^{k}}\geq\sum_{i\in\mathcal{N}}d_{i}\frac{2}{\alpha_{i}^{k+1}}=\frac{2}{c_{\alpha}}\sum_{i\in\mathcal{N}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d_{i}}\tau_{i}^{k}\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{4\rho}{c_{\alpha}}|\mathcal{E}|}\min_{i\in\mathcal{N}}\hat{\tau}_{i}>0,$
[…]
$0\leq$ $t_{k}Q^{k}(\hat{\mathbf{z}})-t_{k+1}Q^{k+1}(\hat{\mathbf{z}})+\mathbf{1}_{\mathcal{I}}(k)\sum_{i\in\mathcal{N}}\Xi_{i}-\frac{\delta^{\prime}}{2}\|\lambda^{k+1}-\lambda^{k}\|^{2}$ (6.28)
@@
which follows from $\mathbf{s}$-update $\mathbf{s}^{k+1}=\mathbf{s}^{k}+\gamma^{k}(1+\eta^{k})\mathbf{x}^{k}-\gamma^{k}\eta^{k}\mathbf{x}^{k-1}$ in (3.11a) and (3.11c). Since $\gamma^{k}=\frac{c_{\gamma}}{\bar{\tau}}\Big(\frac{2}{c_{\alpha}}+\frac{\eta^{k}}{c_{\varsigma}}\Big)^{-1}$ for some $c_{\gamma}\in\Big(0,\frac{1}{2|\mathcal{E}|}\Big)$ and $1\leq\eta^{k}=\max_{i\in\mathcal{N}}\eta_{i}^{k}=\max_{i\in\mathcal{N}}\frac{\tau_{i}^{k-1}}{\tilde{\tau}_{i}^{k}}\leq\max_{i\in\mathcal{N}}\frac{\bar{\tau}_{i}}{\rho\hat{\tau}_{i}}\leq 1/\hat{t}$; hence, $\frac{c_{\gamma}}{\bar{\tau}}\Big(\frac{2}{c_{\alpha}}+\frac{1/\hat{t}}{c_{\varsigma}}\Big)^{-1}\leq\gamma^{k}\leq\frac{c_{\gamma}}{\bar{\tau}}\Big(\frac{2}{c_{\alpha}}+\frac{1}{c_{\varsigma}}\Big)^{-1}$ for all $k\geq 0$. Therefore, $(\lambda^{k_{n}+1}-\lambda^{k_{n}})/\gamma^{k_{n}}\to 0$ and $\eta^{k_{n}}(\mathbf{x}^{k_{n}}-\mathbf{x}^{k_{n}-1})\to 0$ as $n\to\infty$, which together with (6.30) implies that
@@
hence, because $\bar{\tau}_{i}\geq\tau_{i}^{k}\geq\rho\min_{i\in\mathcal{N}}\hat{\tau}_{i}>0$ and $\sigma_{i}^{k}=\zeta_{i}\tau_{i}^{k}$ for some $\zeta_{i}>0$ for all $k\geq 0$, taking the limit on both sides of (6.33) and (6.34), and using [rockafellar1997convex, Theorem 24.4], we get
@@
hence, combining it with (6.32) implies $(\theta^{*},\lambda^{*})\in\operatorname*{\mathbf{argmax}}\{\mathcal{L}(\mathbf{x}^{*},\theta,\lambda):\ \theta\in\mathcal{K}^{*}\cap\mathcal{B},\lambda\}$ and

### v2

@@ 6. Convergence guarantees for D-APDB and the Proof of Main Results
where we used $\eta^{k}(\alpha_{i}^{k}+\beta_{i}^{k})=(c_{\alpha}+c_{\beta})/\tau_{i}^{k}$ for the first inequality, and $\delta\leq 1-(c_{\alpha}+c_{\beta}+c_{\varsigma})$ in the second inequality. For all $i\in\mathcal{N}$, recall that for the case $L_{g_{i}}>0$, we have $\|\theta_{i}^{k}\|\leq B_{i}$; hence, $\|\theta_{i}^{k+1}-\theta_{i}^{k}\|^{2}\leq 4B_{i}^{2}$ for all $k\geq 0$. Therefore, for all $i\in\mathcal{N}$, since $\tau_{i}^{k}\leq\bar{\tau}_{i}$, using $\|x_{i}^{k+1}-x_{i}^{k}\|^{2}\leq 4D_{i}^{2}$ together with $\delta\in(0,1)$, we get $\bar{P}_{i}^{k}\leq-\frac{\delta}{2\tau_{i}^{k}}\|x_{i}^{k+1}-x_{i}^{k}\|^{2}-\frac{\delta}{2\sigma_{i}^{k}}\|\theta_{i}^{k+1}-\theta_{i}^{k}\|^{2}+\Xi^{k}_{i}$, for all $k\in\mathcal{I}$. ∎
@@
Under the premise of Theorem 5.4, given some arbitrary initial points $(x_{i}^{0},\theta_{i}^{0})\in\mathop{\bf dom}\phi_{i}\times\mathcal{K}_{i}^{*}$ and the step size parameters $\bar{\tau}_{i},\zeta_{i}>0$ for $i\in\mathcal{N}$, and the contraction coefficient $\rho\in(0,1)$, it holds for $i\in\mathcal{N}$ that the D-APDB dual iterate sequence $\{\theta_{i}^{k}\}_{k\geq 0}$ is bounded; hence, for $i\in\mathcal{N}$ such that $L_{g_{i}}=0$, there exists $B_{i}>0$ such that $\|\theta_{i}^{k}\|\leq B_{i}$ for all $k\geq 0$.
Proof.
[…]
$\geq\frac{\hat{t}}{2}\sum_{i\in\mathcal{N}}\Big(\frac{\delta}{\bar{\tau}_{i}}\|x_{i}^{k}-\hat{x}_{i}\|^{2}+\frac{1}{\zeta_{i}\bar{\tau}_{i}}\|\theta_{i}^{k}-\hat{\theta}_{i}\|^{2}\Big)+\frac{{2}\hat{t}\rho}{c_{\alpha}}{|\mathcal{E}|}~\min_{i\in\mathcal{N}}\{\hat{\tau}_{i}\}~\|\lambda^{k}-\hat{\lambda}\|^{2}>0,$
$\frac{1}{\gamma^{k}}-\sum_{i\in\mathcal{N}}d_{i}\frac{\eta^{k}}{\varsigma_{i}^{k}}\geq\sum_{i\in\mathcal{N}}d_{i}\frac{2}{\alpha_{i}^{k+1}}=\frac{2}{c_{\alpha}}\sum_{i\in\mathcal{N}}{d_{i}}\tau_{i}^{k}\geq{\frac{4\rho}{c_{\alpha}}|\mathcal{E}|}\min_{i\in\mathcal{N}}\hat{\tau}_{i}>0,$
[…]
$0\leq$ $t_{k}Q^{k}(\hat{\mathbf{z}})-t_{k+1}Q^{k+1}(\hat{\mathbf{z}})+\mathbf{1}_{\mathcal{I}}(k)\sum_{i\in\mathcal{N}}\Xi^{k}_{i}-\frac{\delta^{\prime}}{2}\|\lambda^{k+1}-\lambda^{k}\|^{2}$ (6.4)
[…]
Lemma 6.5.
Under the premise of Lemma 6.3, it holds that
where $\Xi\triangleq\Big(\Big\lfloor\log_{1/\rho}\Big(\max_{i\in\mathcal{N}}\frac{\bar{\tau}_{i}}{\hat{\tau}_{i}}\Big)\Big\rfloor+1\Big)\sum_{i\in\mathcal{N}}\Xi_{i}$ such that for $i\in\mathcal{N}$ with $L_{g_{i}}>0$, $\Xi_{i}\triangleq 2\Big(L_{f_{i}}+\frac{\bar{\tau}_{i}}{c_{\beta}}B_{i}^{2}L_{g_{i}}^{2}\Big)D_{i}^{2}+4\frac{\bar{\tau}_{i}}{c_{\alpha}}C_{g_{i}}^{2}B_{i}^{2}$ using $B_{i}$ given in Assumption 4, and for $i\in\mathcal{N}$ with $L_{g_{i}}=0$, $\Xi_{i}=2L_{f_{i}}D_{i}^{2}+4\frac{\bar{\tau}_{i}}{c_{\alpha}}C_{g_{i}}^{2}B_{i}^{2}$ for $B_{i}>0$ given in Theorem 6.4.
Proof.
The desired results immediately follows from the bound on the cardinality of $\mathcal{I}$ given in Lemma 6.2 and the fact that $t_{k}\leq t_{0}=1$ for all $k\geq 0$. ∎
Theorem 6.6.
Under the premise of Lemma 6.3. Let $\{x_{i}^{k},\theta_{i}^{k}\}_{k\geq 0}$ be the D-APDB iterate sequence for $i\in\mathcal{N}$, and $\{s_{i}^{k}\}_{k\geq 0}$ denote the D-APDB auxiliary sequence with $s_{i}^{0}=0$ for $i\in\mathcal{N}$, and define $\{\lambda^{k}\}_{k\geq 0}$ such that $\lambda^{k}=A\mathbf{s}^{k}$ with $\mathbf{s}^{k}=[s_{i}^{k}]_{i\in\mathcal{N}}$ for all $k\geq 0$. For any $\mathbf{x}\in\mathop{\bf dom}\phi$, $\theta\in\mathcal{K}^{*}\cap\mathcal{B}$ and $\lambda\in\mathbb{R}^{n|\mathcal{E}|}$, let $\mathbf{z}\triangleq(\mathbf{x},\mathbf{y})$ with $\mathbf{y}\triangleq(\theta,\lambda)$. Then, for any $K\geq 1$, it holds that
@@
holds for any $K\geq 1$, where $T_{K}=\sum_{k=0}^{K-1}t_{k}$ and $(\bar{\mathbf{x}}^{K},\bar{\mathbf{y}}^{K})\triangleq\frac{1}{T_{K}}\sum_{k=0}^{K-1}t_{k}(\mathbf{x}^{k},\mathbf{y}^{k})$. Therefore, Lemma 4.4 and Lemma 6.3 together imply that
@@
Using the same arguments we employed while deriving (4.3) and (4.12), we can further lower bound the left hand side of (6.8) as below:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Gamma^{K}(\mathbf{z})}$ $\geq\mathcal{L}(\bar{\mathbf{x}}^{K},\mathbf{y})-\mathcal{L}(\mathbf{x},\bar{\mathbf{y}}^{K})+\frac{1}{2\gamma^{K-1}}\frac{t_{K-1}}{T_{K}}\|\lambda-\lambda^{K}\|^{2}+\frac{t_{K-1}}{2T_{K}}\sum_{i\in\mathcal{N}}\varsigma_{i}^{K}\|x_{i}^{K}-x_{i}^{K-1}\|^{2}$
@@
Let $(x^{*},\theta^{*})\in\mathbb{R}^{n}\times\mathcal{K}^{*}$ denote an arbitrary primal-dual optimal pair for (1.1) as defined in Assumption 4 such that $\|\theta_{i}^{*}\|\leq B_{i}/2$ for $i\in\mathcal{N}$ with $L_{g_{i}}>0$, and $\lambda^{*}\in\mathbb{R}^{n|\mathcal{E}|}$ be an optimal dual variable corresponding to the consensus constraint $A\mathbf{x}=0$ in (2.1). Under the premise of Lemma 6.3, initialized from arbitrary $x_{i}^{0}\in\mathop{\bf dom}\phi_{i}$ and $\theta_{i}^{0}=0$ for all $i\in\mathcal{N}$, the ergodic iterate sequences of D-APDB, i.e., $\{(\bar{x}_{i}^{k},\bar{\theta}_{i}^{k})\}_{k\geq 0}$ for all $i\in\mathcal{N}$, satisfy
@@
$\mathcal{L}(\bar{\mathbf{x}}^{K},\theta^{*},\lambda^{*})-\mathcal{L}(\mathbf{x}^{*},\bar{\theta}^{K},\bar{\lambda}^{K})\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\Gamma}(\mathbf{x}^{*},\mathbf{y}^{*})}/T_{K}.$
@@
which also satisfy $\left\langle\tilde{w}_{i}^{(1)},~\tilde{w}_{i}^{(2)}\right\rangle=0$ for $i\in\mathcal{N}$. Thus, for all $i\in\mathcal{N}$, $\|\tilde{w}_{i}^{(2)}\|=\|\mathcal{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\mathcal{K}_{i}}}(\tilde{w}_{i})-\tilde{w}_{i}\|\triangleq d_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\mathcal{K}_{i}}}(\tilde{w}_{i}).$ Next, we define $\tilde{\theta}=[\tilde{\theta}_{i}]_{i\in\mathcal{N}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\in\mathcal{K}^{*}\cap\mathcal{B}}$ such that
$\tilde{\theta}_{i}\triangleq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\|\theta_{i}^{*}\|}\cdot\frac{\tilde{w}_{i}^{(2)}}{\|\tilde{w}_{i}^{(2)}\|}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{K}_{i}^{*}\cap\mathcal{B}_{i}},\quad\forall~i\in\mathcal{N}.$
@@
Similarly, we define $\tilde{\lambda}\triangleq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2\|\lambda^{*}\|}\cdot\frac{A\bar{\mathbf{x}}^{K}}{\|A\bar{\mathbf{x}}^{K}\|}$; hence, from the definition, we get
@@
Note that $\mathcal{L}(\mathbf{x}^{*},\bar{\theta}^{K},\bar{\lambda}^{K})\leq\mathcal{L}(\mathbf{x}^{*},\theta^{*},\lambda^{*})=\varphi^{*}$ since $(\mathbf{x}^{*},\theta^{*},\lambda^{*})$ is a saddle point of $\mathcal{L}$, and that $\tilde{\theta}\in\mathcal{K}^{*}\cap\mathcal{B}$. Thus (6.12) and (6.13) imply
@@
Moreover, since $\mathbf{x}^{*}\in\mathop{\bf dom}\phi$ and $\tilde{\theta}\in\mathcal{K}^{*}\cap\mathcal{B}$, (6.6) also implies that
@@
Moreover, using the saddle point property of $(\mathbf{x}^{*},\theta^{*},\lambda^{*})$ one more time, we see that
$\mathcal{L}(\bar{\mathbf{x}}^{K},{\theta}^{*},{\lambda}^{*})-\mathcal{L}(\mathbf{x}^{*},\theta^{*},\lambda^{*})\geq 0.$
It follows that
@@
Recall that $g_{i}(\bar{x}_{i}^{K})=\tilde{w}_{i}$ such that (6.11) holds for $i\in\mathcal{N}$. Hence, we immediately obtain
$\langle\theta_{i}^{*},g_{i}(\bar{x}_{i}^{K})\rangle=\langle\theta_{i}^{*},~\tilde{w}_{i}-\tilde{w}_{i}^{(1)}+\tilde{w}_{i}^{(1)}\rangle\leq\langle\theta_{i}^{*},~\tilde{w}_{i}-\tilde{w}_{i}^{(1)}\rangle\leq\|\theta_{i}^{*}\|d_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-\mathcal{K}_{i}}}(g_{i}(\bar{x}_{i}^{K})).$
Thus (6.18) implies
@@
Combining (6.17) and (6.19) immediately yields the desired result. ∎
@@
Under the premise of Theorem 5.4, given some arbitrary initial points $(x_{i}^{0},\theta_{i}^{0})\in\mathop{\bf dom}\phi_{i}\times\mathcal{K}_{i}^{*}$ and the step size parameters $\bar{\tau}_{i},\zeta_{i}>0$ for $i\in\mathcal{N}$, and the contraction coefficient $\rho\in(0,1)$, let $\{x_{i}^{k},\theta_{i}^{k}\}_{k\geq 0}$ denote the D-APDB iterate sequence for $i\in\mathcal{N}$. Then, there exists $(x^{*},\theta^{*})$ a primal-dual optimal solution to (1.1) such that $\{(x_{i}^{k},\theta_{i}^{k})\}_{k\geq 0}$ converges to $(x^{*},\theta_{i}^{*})$ for all $i\in\mathcal{N}$, i.e., $\lim_{k\to\infty}x_{i}^{k}=x^{*}$ and $\lim_{k\to\infty}\theta_{i}^{k}=\theta_{i}^{*}$ for $i\in\mathcal{N}$.
@@
which follows from $\mathbf{s}$-update $\mathbf{s}^{k+1}=\mathbf{s}^{k}+\gamma^{k}(1+\eta^{k})\mathbf{x}^{k}-\gamma^{k}\eta^{k}\mathbf{x}^{k-1}$ in (3.9a) and (3.9c). Since $\gamma^{k}=\frac{c_{\gamma}}{\bar{\tau}}\Big(\frac{2}{c_{\alpha}}+\frac{\eta^{k}}{c_{\varsigma}}\Big)^{-1}$ for some $c_{\gamma}\in\Big(0,\frac{1}{2|\mathcal{E}|}\Big)$ and $1\leq\eta^{k}=\max_{i\in\mathcal{N}}\eta_{i}^{k}=\max_{i\in\mathcal{N}}\frac{\tau_{i}^{k-1}}{\tilde{\tau}_{i}^{k}}\leq\max_{i\in\mathcal{N}}\frac{\bar{\tau}_{i}}{\rho\hat{\tau}_{i}}\leq 1/\hat{t}$; hence, $\frac{c_{\gamma}}{\bar{\tau}}\Big(\frac{2}{c_{\alpha}}+\frac{1/\hat{t}}{c_{\varsigma}}\Big)^{-1}\leq\gamma^{k}\leq\frac{c_{\gamma}}{\bar{\tau}}\Big(\frac{2}{c_{\alpha}}+\frac{1}{c_{\varsigma}}\Big)^{-1}$ for all $k\geq 0$. Therefore, $(\lambda^{k_{n}+1}-\lambda^{k_{n}})/\gamma^{k_{n}}\to 0$ and $\eta^{k_{n}}(\mathbf{x}^{k_{n}}-\mathbf{x}^{k_{n}-1})\to 0$ as $n\to\infty$. Thus, (6.23) implies that
@@
hence, because $\bar{\tau}_{i}\geq\tau_{i}^{k}\geq\rho\min_{i\in\mathcal{N}}\hat{\tau}_{i}>0$ and $\sigma_{i}^{k}=\zeta_{i}\tau_{i}^{k}$ for some $\zeta_{i}>0$ for all $k\geq 0$, taking the limit on both sides of (6.26) and (6.27), and using [45, Theorem 24.4], we get
@@
hence, (6.25) implies $(\theta^{*},\lambda^{*})\in\operatorname*{\mathbf{argmax}}\{\mathcal{L}(\mathbf{x}^{*},\theta,\lambda):\ \theta\in\mathcal{K}^{*}\cap\mathcal{B},\lambda\}$ and